OpenStudy (anonymous):

Plz help!! solve each of the polynomial equations x^3 - 7x^2 -x +7=0

OpenStudy (misty1212):

HI!!

OpenStudy (dayakar):

sum of the coefficients all powers of x= 1-7-1+7 = 8-8 =0

OpenStudy (dayakar):

therefore x-1 is a factor of above equation

OpenStudy (dayakar):

r u there

OpenStudy (dayakar):

sum of coefficients of odd powers of x = -1+1 =0 ------(I) sum of coefficients of even powers of x = -7+7 =0------(II) (I)= (II) therefore ,x+1 is a factor of above equation

OpenStudy (dayakar):

x-7 is another factor

OpenStudy (dayakar):

(x-1)(x+1)(x-7)=0